OpenStudy (anonymous):

what is the concentration of 42.45 ml of hcl when titrated with 25ml of 0.1 moldm3 of naoh??

OpenStudy (aaronq):

at the equivalence point the moles of base are equal to the moles of acid, right? \(Molarity=\dfrac{n_{solute}}{L_{solution}}\rightarrow n_{solute}=M*L_{solution}\) \(n_{acid}=n_{base}\) so, by substitution: \(M_{acid}*L_{solution}=M_{base}*L_{solution}\) which is the \(M_1V_1=M_2V_2\) equation.

OpenStudy (anonymous):

what exactly does the n and l stand for??

OpenStudy (aaronq):

n is moles and L is liters

OpenStudy (anonymous):

can you also please show me how to get the no. of moles of NaOH?

OpenStudy (aaronq):

use the molarity formula (above), plug in your values for NaOH, solve for n.

OpenStudy (anonymous):

would it be 2.5 moles? sorry to ask so many questions

OpenStudy (aaronq):

it would be n=M*L=(0.1 mol/L)*(0.025 L)=0.0025 moles

OpenStudy (aaronq):

remember that the volume is in terms of Liters, so you will need to convert.

OpenStudy (anonymous):

so that means that the no. of moles of hcl is also 0.0025?

OpenStudy (aaronq):

indeed

OpenStudy (anonymous):

is the concentration 0.0589?

OpenStudy (aaronq):

thats right. don't forget the units though

OpenStudy (anonymous):

\[moldm ^{3}\]

OpenStudy (aaronq):

it should be "mol/\(dm^3\)", because it's moles per liter.

OpenStudy (anonymous):

thank you so much :)

OpenStudy (aaronq):

no problem!